a light rope connects two blocks with mass of 5kgs and 10kgs as shown. find the acceleration of the system and the tensions in the cable.
1
Expert's answer
2021-08-04T14:15:38-0400

Solution.

"m_1=5kg;"

"m_2=10kg;"

"m_1a=T-m_1g;"

"m_2a=m_2g-T;"

"m_1a+m_2a=m_2g-m_1g;"

"a(m_1+m_2)=m_2g-m_1g;"

"a=\dfrac{m_2g-m_1g}{m_1+m_2};"

"a=\dfrac{10\sdot9.8-5\sdot9.8}{5+10}=3.27m/s^2;"

"T=m_1a+m_1g;"

"T=5\sdot9.8+5\sdot3.27=65.35N;"

Answer: "a=3.27m\/s^2; T=65.35N."